### v1

# arXiv:2603.03228v1[math.GN] 03 Mar 2026

Abstract
We construct the Extended Real Line with Reentry (ERI), a quotient of $\overline{\mathbb{R}}=[-\infty,+\infty]$ obtained by collapsing $\{-\infty,0,+\infty\}$ to a single point $\ast$ and imposing a density condition on neighborhoods of $\ast$. ERI is compact, $T_{1}$, path-connected, US, and not KC, providing an explicit, constructive example separating US from KC in the Wilansky hierarchy. We give a complete convergence criterion at $\ast$, identify the failure of first-countability as the mechanism enabling US without Hausdorff separation, and locate ERI in the refined hierarchy of Clontz [2]: ERI is SC (sequentially closed) but not weakly Hausdorff. The construction generalizes to arbitrary compact Hausdorff spaces without isolated points, and the only continuous real-valued functions on ERI are the constants.
MSC 2020: 54D10, 54D25, 54D30, 54G20. Keywords: US space, KC space, Wilansky hierarchy, compact quotient, density condition, sequential convergence, weakly Hausdorff, sequentially closed.
1 Introduction
1.1 The Wilansky hierarchy
The study of separation axioms strictly between $T_{1}$ and $T_{2}$ (Hausdorff) was initiated by Wilansky [1], who introduced the following properties for a topological space $X$:
Definition 1.1 (Wilansky [1]).
  1. (i)
    $X$ is a US-space (uniquely sequential) if every convergent sequence in $X$ has a unique limit.
  2. (ii)
    $X$ is a KC-space (kompacts closed) if every compact subset of $X$ is closed.
The implications
$T_{2}\;\Longrightarrow\;\mathrm{KC}\;\Longrightarrow\;\mathrm{US}\;\Longrightarrow\;T_{1}$ (1)
are all strict [1]. A refined hierarchy with additional intermediate axioms (weakly Hausdorff, $k$-Hausdorff, semi-Hausdorff) has been developed by Clontz [2]. The role of first-countability is clarified by the classical equivalence:
Fact 1.2 (Folklore; see e.g. [5, p. 60]).
A first-countable space is Hausdorff if and only if it is US.
Thus in first-countable spaces, the four levels of (1) collapse to two ($T_{1}$ and $T_{2}$), and only in the absence of first-countability can US and $T_{2}$ genuinely separate.
1.2 Known examples
Several examples illustrate the separation between US and $T_{2}$:
  • The cocountable topology on an uncountable set is $T_{1}$ and KC (hence US), but neither $T_{2}$ nor compact (it is Lindelöf; see [9], Ex. 20). Its sequential structure is trivial: the only convergent sequences are the eventually constant ones.
  • Van Douwen’s example [3] is a countable, compact, Fréchet–Urysohn, US space that is anti-Hausdorff. By Künzi and van der Zypen [4], every Fréchet US-space is KC, so van Douwen’s space is also KC. The construction uses maximal almost disjoint (MAD) families and is highly non-constructive.
  • The line with two origins (LTO) is $T_{1}$ and first-countable but not US: the sequence $1/n$ converges to both origins.
Three compact US-not-KC examples are recorded in pi-Base [12]:
  • The product $\mathbb{Q}^{\ast}\times\mathbb{Q}^{\ast}$ of two copies of the one-point compactification of the rationals (pi-Base S31); it is weakly Hausdorff but not KC [2].
  • The ordinal space $\omega_{1}+1$ with doubled endpoint (pi-Base S37); US but not $k_{2}$-Hausdorff.
  • The one-point compactification of the Arens–Fort space (pi-Base S165); US but not weakly Hausdorff.
None of these is path-connected or arises from an explicit elementary construction with a transparent convergence criterion.
1.3 Contribution
We introduce the Extended Real Line with Reentry (ERI), a compact quotient space arising from an elementary density condition, and prove:
  1. (a)
    ERI is US (Theorem 5.1);
  2. (b)
    ERI is not KC (Theorem 5.2);
  3. (c)
    ERI is not first-countable at $\ast$ (Proposition 4.1), and this failure is the precise mechanism enabling US without $T_{2}$ (Corollary 5.5).
While compact US-not-KC spaces are known (see §1.2 above), ERI is distinguished by the combination of:
  • Path-connectedness: the previously known compact US-not-KC examples (products, ordinal spaces, Arens–Fort compactifications) are totally disconnected or not path-connected.
  • Constructive and explicit: unlike van Douwen’s MAD-family construction, ordinal constructions using $\omega_{1}$, or the Arens–Fort space.
  • Complete convergence criterion: ERI has a precise characterization of sequential convergence (Proposition 4.2).
  • Transparent mechanism: a single density condition on neighborhoods of $\ast$ is responsible for all non-Hausdorff phenomena.
The paper is organized as follows. Section 2 defines the ERI space and establishes its basic structural properties. Section 3 proves the separation properties ($T_{1}$, not $T_{2}$) and the topological properties (connectedness, path-connectedness, compactness). Section 4 studies first-countability and sequential convergence. Section 5 contains the main results: ERI is US but not KC, and the role of first-countability is identified. Section 6 compares ERI with known examples. Section 7 generalizes the construction. Section 9 collects concluding remarks.
2 The ERI Construction
2.1 Definition and notation
Definition 2.1 (ERI space).
Let $\overline{\mathbb{R}}=[-\infty,+\infty]$ denote the extended real line with its standard order topology. Define the equivalence relation $\sim$ on $\overline{\mathbb{R}}$ whose sole non-singleton class is $\{-\infty,0,+\infty\}$, and let $X_{0}:=\overline{\mathbb{R}}/{\sim}$ be the quotient set. Write $q:\overline{\mathbb{R}}\to X_{0}$ for the projection, $\ast:=q(-\infty)=q(0)=q(+\infty)$ for the collapsed point, and $\hat{x}:=q(x)$ for all $x\in\overline{\mathbb{R}}$. The map $x\mapsto\hat{x}$ is a bijection from $\mathbb{R}\setminus\{0\}$ onto $X_{0}\setminus\{\ast\}$.
The ERI topology on $X_{0}$ is
[…]
We write $X:=(X_{0},\tau_{\mathrm{ERI}})$.
Remark 2.2 (Standing notation).
Throughout this paper: $\overline{\mathbb{R}}=[-\infty,+\infty]$ denotes the extended real line (compact, Hausdorff, homeomorphic to [1,1][-1,1]); $\mathbb{R}=(-\infty,+\infty)$ the usual real line. A subset $D$ of a topological space $Y$ is dense if $\mathrm{cl}(D)=Y$. A set $A$ is nowhere dense if $\mathrm{int}(\overline{A})=\varnothing$; for closed $A$ this reduces to $\mathrm{int}(A)=\varnothing$. When we say that $F\subset\mathbb{R}\setminus\{0\}$ is closed, we mean closed as a subset of $\overline{\mathbb{R}}$.
Remark 2.3.
$\tau_{\mathrm{ERI}}$ is not the standard quotient topology $\tau_{q}:=\{U\subseteq X_{0}:q^{-1}(U)\text{ is open in }\overline{\mathbb{R}}\}$. Condition (b) imposes an extra density requirement, making $\tau_{\mathrm{ERI}}$ strictly coarser than $\tau_{q}$ (Proposition 2.13).
2.2 Preliminary facts
Lemma 2.4 (No isolated points).
$\overline{\mathbb{R}}$ has no isolated points.
Proof.
Every basic open set in $\overline{\mathbb{R}}$ is an open interval containing infinitely many points. ∎
Corollary 2.5.
For any finite set $A\subset\overline{\mathbb{R}}$, the set $\overline{\mathbb{R}}\setminus A$ is dense in $\overline{\mathbb{R}}$.
Proof.
Each $\overline{\mathbb{R}}\setminus\{a\}$ is dense and open (no isolated points). A finite intersection of dense open sets is dense: if $D_{1},D_{2}$ are dense open and $V$ is non-empty open, then $D_{1}\cap V$ is non-empty and open (a dense set meets every non-empty open set), so $D_{2}\cap(D_{1}\cap V)\neq\varnothing$; iterate. ∎
Definition 2.6 (Saturated set).
$S\subseteq\overline{\mathbb{R}}$ is saturated if $q^{-1}(q(S))=S$. Equivalently, $S\cap\{-\infty,0,+\infty\}\in\{\varnothing,\;\{-\infty,0,+\infty\}\}$.
Lemma 2.7.
  1. (i)
    Every subset of $\mathbb{R}\setminus\{0\}$ is saturated, and its complement in $\overline{\mathbb{R}}$ is also saturated.
  2. (ii)
    If $S$ is saturated, then $q^{-1}(q(S))=S$ and $q(S)^{c}=q(S^{c})$.
Proof.
Immediate from the fact that $\{-\infty,0,+\infty\}$ is the unique non-singleton class. ∎
Fact 2.8.
For any function $f\!:A\to B$ and subsets $U,V\subseteq B$: $f^{-1}(U\cap V)=f^{-1}(U)\cap f^{-1}(V)$. In particular, $f^{-1}(U)\cap f^{-1}(V)\neq\varnothing$ implies $U\cap V\neq\varnothing$.
2.3 Density facts
Lemma 2.9 (Dense meets open).
In any topological space, if $D$ is dense and $V$ is non-empty and open, then $D\cap V\neq\varnothing$. (Standard; see [10], 1.3.5.)
Remark 2.10.
The intersection of two dense open sets in any topological space is dense: if $D_{1},D_{2}$ are dense and open and $V$ is non-empty open, then $D_{1}\cap V$ is non-empty open (by Lemma 2.9), so $D_{2}\cap(D_{1}\cap V)\neq\varnothing$. By induction, finite intersections of dense open sets are dense. The Baire Category Theorem extends this to countable intersections in compact Hausdorff spaces; see [7], §48.
2.4 Topology axioms and basic properties
Proposition 2.11 ($\tau_{\mathrm{ERI}}$ is a topology).
The collection $\tau_{\mathrm{ERI}}$ satisfies the topology axioms.
Proof.
(i) $q^{-1}(\varnothing)=\varnothing$ is open; condition (b) is vacuous. $q^{-1}(X_{0})=\overline{\mathbb{R}}$ is open and dense.
(ii) Arbitrary unions. If $\{U_{\alpha}\}\subseteq\tau_{\mathrm{ERI}}$ and $U=\bigcup_{\alpha}U_{\alpha}$, then $q^{-1}(U)$ is an open union. If $\ast\in U$, some $q^{-1}(U_{\beta})$ is dense, and $q^{-1}(U_{\beta})\subseteq q^{-1}(U)$.
(iii) Finite intersections. If $G_{1},G_{2}\in\tau_{\mathrm{ERI}}$, then $q^{-1}(G_{1}\cap G_{2})=q^{-1}(G_{1})\cap q^{-1}(G_{2})$ is open. If $\ast\in G_{1}\cap G_{2}$, both preimages are dense and open; their intersection is dense by Remark 2.10. ∎
Proposition 2.12 (Continuity).
$q:\overline{\mathbb{R}}\to(X,\tau_{\mathrm{ERI}})$ is continuous.
Proof.
Condition (a) states $q^{-1}(U)$ is open for every $U\in\tau_{\mathrm{ERI}}$. ∎
Proposition 2.13 (Strict inclusion).
$\tau_{\mathrm{ERI}}\subsetneq\tau_{q}$. The standard quotient $(X_{0},\tau_{q})$ is Hausdorff.
Proof.
The inclusion $\tau_{\mathrm{ERI}}\subseteq\tau_{q}$ is immediate. For strictness, given $x\in\mathbb{R}\setminus\{0\}$, the set $V_{\ast}:=q([-\infty,x-\varepsilon)\cup(x+\varepsilon,+\infty])$ with $\varepsilon=|x|/2$ belongs to $\tau_{q}$ (preimage is open and contains $\{-\infty,0,+\infty\}$) but not to $\tau_{\mathrm{ERI}}$ (the preimage has complement $x-\varepsilon,x+\varepsilon]$ with non-empty interior, so it is not dense). The Hausdorff property of $\tau_{q}$ follows by explicit $\varepsilon$-separation: given distinct $\hat{a},\hat{b}\in X_{0}$, their preimages under $q$ are singletons or $\{-\infty,0,+\infty\}$, all of which can be separated by disjoint open intervals in $\overline{\mathbb{R}}$. ∎
2.5 Neighborhoods of $\ast$
Definition 2.14 (The family $\mathcal{W}$).
For any closed nowhere-dense $F\subset\mathbb{R}\setminus\{0\}$, define $W_{F}:=q(\overline{\mathbb{R}}\setminus F)$.
Lemma 2.15 (Canonical neighborhoods).
  1. (i)
    If $U\in\tau_{\mathrm{ERI}}$ and $\ast\in U$, then $F_{U}:=\overline{\mathbb{R}}\setminus q^{-1}(U)$ is a closed nowhere-dense subset of $\mathbb{R}\setminus\{0\}$, and $U=W_{F_{U}}$.
  2. (ii)
    For every closed nowhere-dense $F\subset\mathbb{R}\setminus\{0\}$, $W_{F}$ is an open neighborhood of $\ast$ in $\tau_{\mathrm{ERI}}$, with $q^{-1}(W_{F})=\overline{\mathbb{R}}\setminus F$.
Proof.
(i).  Set $V:=q^{-1}(U)$. Conditions (a) and (b) give $V$ open and dense. Since $\{-\infty,0,+\infty\}\subseteq V$, we get $F_{U}\subseteq\mathbb{R}\setminus\{0\}$. Then $F_{U}$ is closed ($V$ is open) with $\mathrm{int}(F_{U})=\varnothing$ ($V$ is dense), hence nowhere dense. Since $F_{U}\subset\mathbb{R}\setminus\{0\}$, $\overline{\mathbb{R}}\setminus F_{U}$ is saturated (Lemma 2.7(i)), giving $U=W_{F_{U}}$.
(ii).  $\overline{\mathbb{R}}\setminus F$ contains $\{-\infty,0,+\infty\}$, is saturated, open ($F$ closed), and dense ($F$ nowhere dense). Both conditions hold; $\ast\in W_{F}$. ∎
Remark 2.16.
Every closed subset of $\overline{\mathbb{R}}$ contained in $\mathbb{R}\setminus\{0\}$ is compact (closed in a compact space) and contained in $M,-\delta]\cup[\delta,M]$ for some $M,\delta>0$.
2.6 The key intersection principle
Proposition 2.17 (Key intersection principle (KIP)).
If $U\in\tau_{\mathrm{ERI}}$ with $\ast\in U$ and $V\in\tau_{\mathrm{ERI}}$ with $V\neq\varnothing$, then $U\cap V\neq\varnothing$. In particular, every open neighborhood of $\ast$ is dense in $X$.
Proof.
$q^{-1}(U)$ is dense (condition (b)) and $q^{-1}(V)$ is non-empty open (condition (a)). By Lemma 2.9, $q^{-1}(U)\cap q^{-1}(V)\neq\varnothing$; by Fact 2.8, $U\cap V\neq\varnothing$. ∎
Corollary 2.18.
$\ast\in\mathrm{cl}(O)$ for every non-empty open $O\subseteq X$.
Proof.
Every open neighborhood $W$ of $\ast$ satisfies $W\cap O\neq\varnothing$ by the KIP. ∎
3 Separation and Topological Properties
Proposition 3.1 ($X$ is $T_{1}$).
Every singleton in $X$ is closed.
Proof.
Case $y=\ast$: $q^{-1}(X\setminus\{\ast\})=\overline{\mathbb{R}}\setminus\{-\infty,0,+\infty\}$ is open (finite sets are closed in the Hausdorff $\overline{\mathbb{R}}$). Since $\ast\notin X\setminus\{\ast\}$, condition (b) is vacuous.
Case $y=\hat{x}$, $x\in\mathbb{R}\setminus\{0\}$: $q^{-1}(X\setminus\{\hat{x}\})=\overline{\mathbb{R}}\setminus\{x\}$ is open and dense (Lemma 2.4). Since $\ast\in X\setminus\{\hat{x}\}$, condition (b) is satisfied. ∎
Theorem 3.2 (Failure of Hausdorff).
$(X,\tau_{\mathrm{ERI}})$ is not Hausdorff: $\ast$ cannot be separated from any other point by disjoint open sets.
Proof.
Let $U,V\in\tau_{\mathrm{ERI}}$ with $\ast\in U$ and $\hat{x}\in V$. Since $V\neq\varnothing$, the KIP (Proposition 2.17) gives $U\cap V\neq\varnothing$. ∎
Proposition 3.3 (Compactness).
$X$ is compact.
Proof.
$(X_{0},\tau_{q})$ is compact (continuous image of compact $\overline{\mathbb{R}}$) and $\tau_{\mathrm{ERI}}\subseteq\tau_{q}$. ∎
Proposition 3.4 (Connectedness).
$X$ is connected.
Proof.
If $U,V$ partition $X$ with $\ast\in U$, then $V$ is non-empty open and the KIP gives $U\cap V\neq\varnothing$, a contradiction. ∎
Proposition 3.5 (Path-connectedness).
$X$ is path-connected.
Proof.
Given $\hat{a},\hat{b}\in X$, choose representatives $a,b\in\mathbb{R}$ (using 0 for $\ast$). The linear path $\tilde{\gamma}(t)=(1-t)a+tb$ lies in $\mathbb{R}\subset\overline{\mathbb{R}}$ and is continuous. Then $\gamma:=q\circ\tilde{\gamma}$ is a continuous path from $\hat{a}$ to $\hat{b}$. ∎
4 Countability and Sequential Properties
4.1 First countability fails at $\ast$
Proposition 4.1 (Not first-countable at $\ast$).
$\ast$ has no countable neighborhood base in $X$. In particular, $X$ is neither first-countable nor second-countable.
Proof.
Suppose $\{U_{n}\}_{n\geq 1}$ is a countable neighborhood base at $\ast$. By Lemma 2.15(i), $U_{n}=W_{F_{n}}$ for closed nowhere-dense $F_{n}$, and $V_{n}:=q^{-1}(U_{n})=\overline{\mathbb{R}}\setminus F_{n}$ is dense and open.
Since $\overline{\mathbb{R}}$ is compact Hausdorff, it is a Baire space ([7], §48). Hence $D:=\bigcap_{n\geq 1}V_{n}$ is dense. Pick $t\in D\cap(\mathbb{R}\setminus\{0\})$ (using that $(-1,0)\cup(0,1)$ is non-empty open). Then $W_{\{t\}}$ is an open neighborhood of $\ast$ (Lemma 2.15(ii)), so some $U_{n_{0}}\subseteq W_{\{t\}}$, giving $q^{-1}(U_{n_{0}})\subseteq\overline{\mathbb{R}}\setminus\{t\}$. But $t\in D\subseteq V_{n_{0}}=q^{-1}(U_{n_{0}})$, a contradiction. ∎
4.2 Sequential convergence
Throughout this subsection, $(x_{n})$ denotes a sequence in $\mathbb{R}\setminus\{0\}$ and $\hat{x}_{n}:=q(x_{n})\in X$.
Proposition 4.2 (Convergence criterion).
$\hat{x}_{n}\to\ast$ in $X$ if and only if $(x_{n})$ has no accumulation point in $\mathbb{R}\setminus\{0\}$.
Proof.
($\Rightarrow$).  Suppose $t\in\mathbb{R}\setminus\{0\}$ is an accumulation point. A subsequence satisfies $x_{n_{k}}\to t$. Let $C:=\{x_{n_{k}}:k\geq 1\}\cup\{t\}$. Since $x_{n_{k}}\to t$, the only accumulation point of $C$ in $\overline{\mathbb{R}}$ is $t$, and $t\in C$, so $C$ is closed. $C$ is countable, hence nowhere dense ($\overline{\mathbb{R}}\cong[-1,1]$, so every non-empty open set is uncountable). Since $C\subset\mathbb{R}\setminus\{0\}$, $W_{C}=q(\overline{\mathbb{R}}\setminus C)\in\tau_{\mathrm{ERI}}$ contains $\ast$ (Lemma 2.15(ii)). But $\hat{x}_{n_{k}}\notin W_{C}$ for all $k$. If $\hat{x}_{n}\to\ast$, then $(\hat{x}_{n})$ would be eventually in $W_{C}$, forcing $\hat{x}_{n_{k}}\in W_{C}$ for large kk—a contradiction.
($\Leftarrow$).  Let $U$ be an open neighborhood of $\ast$. By Lemma 2.15(i), $F:=\overline{\mathbb{R}}\setminus q^{-1}(U)$ is a compact subset of $\mathbb{R}\setminus\{0\}$ (Remark 2.16). If infinitely many $x_{n}$ belonged to $F$, the Bolzano–Weierstrass theorem yields a convergent subsequence with limit in $F\subset\mathbb{R}\setminus\{0\}$, contradicting the hypothesis. Hence eventually $\hat{x}_{n}\in U$. ∎
Remark 4.3 (Equivalent formulations).
The condition $(x_{n})$ has no accumulation point in $\mathbb{R}\setminus\{0\}$” is equivalent to:
  1. (a)
    every accumulation point of $(x_{n})$ in $\overline{\mathbb{R}}$ belongs to $\{-\infty,0,+\infty\}$;
  2. (b)
    $(x_{n})$ eventually leaves every compact subset of $\mathbb{R}\setminus\{0\}$.
In particular, the sequence need not converge in $\overline{\mathbb{R}}$: different subsequences may tend to different elements of $\{-\infty,0,+\infty\}$.
4.3 Uniqueness of sequential limits
Proposition 4.4 (Unique sequential limits).
Limits of convergent sequences in $X$ are unique.
Proof.
Let $(y_{n})$ be a sequence in $X$ converging to $L_{1}$ and $L_{2}$ with $L_{1}\neq L_{2}$.
Reduction to sequences in $X\setminus\{\ast\}$.  If $y_{n}=\ast$ for infinitely many $n$, then every open $U\ni L$ must contain $\ast$ for any limit $L$. Since $\{\ast\}$ is closed (Proposition 3.1), $X\setminus\{\ast\}$ is an open set containing $L$ but not $\ast$ whenever $L\neq\ast$—a contradiction. Hence $L=\ast$ is the only possible limit. If $y_{n}=\ast$ for only finitely many $n$, the tail lies in $X\setminus\{\ast\}$. In either case we reduce to $\hat{x}_{n}$ with $x_{n}\in\mathbb{R}\setminus\{0\}$.
Case 1: $\hat{x}_{n}\to\ast$ and $\hat{x}_{n}\to\hat{c}$, $c\in\mathbb{R}\setminus\{0\}$.  For small $\varepsilon>0$, $q((c-\varepsilon,c+\varepsilon))$ is an open neighborhood of $\hat{c}$ not containing $\ast$. Convergence to $\hat{c}$ forces $x_{n}$ eventually into $(c-\varepsilon,c+\varepsilon)$, making $c$ an accumulation point in $\mathbb{R}\setminus\{0\}$—contradicting Proposition 4.2.
Case 2: $\hat{x}_{n}\to\hat{c}$ and $\hat{x}_{n}\to\hat{d}$, $c\neq d$ both in $\mathbb{R}\setminus\{0\}$.  Disjoint neighborhoods $q((c-\varepsilon,c+\varepsilon))$ and $q((d-\varepsilon,d+\varepsilon))$ cannot both eventually contain $\hat{x}_{n}$. ∎
4.4 Nets witness non-Hausdorff
Example 4.5 (A net with two limits).
Fix $c\in\mathbb{R}\setminus\{0\}$. Let $\mathcal{D}:=\mathcal{N}_{\ast}\times\mathcal{N}_{\hat{c}}$ (products of neighborhood filters), directed by reverse inclusion. For each $(U,V)\in\mathcal{D}$, the KIP gives $U\cap V\neq\varnothing$; select $z_{(U,V)}\in U\cap V$ (Axiom of Choice). Then $z_{(U,V)}\to\ast$ (given $W\ni\ast$, $(U,V)\succeq(W,X_{0})$ implies $z_{(U,V)}\in U\subseteq W$) and $z_{(U,V)}\to\hat{c}$ (similarly).
Remark 4.6.
The $T_{1}$ axiom alone does not guarantee unique sequential limits. In the line with two origins, $1/n$ converges to both origins. The key ingredient in Proposition 4.4 is the convergence criterion (Proposition 4.2), specific to the ERI topology.
4.5 Non-sequentiality
Recall that a topological space is sequential (Franklin [6]) if every sequentially closed set is closed; equivalently, if the topology is completely determined by convergent sequences.
Proposition 4.7.
$(X,\tau_{\mathrm{ERI}})$ is not a sequential space.
Proof.
The set $K=q([1,2])$ is sequentially closed but not closed. It is not closed by Theorem 5.2. To see that $K$ is sequentially closed, suppose $\hat{x}_{n}\in K$ (i.e., $x_{n}\in[1,2]$) and $\hat{x}_{n}\to L$ in $X$.
  • If $L=\hat{c}$ with $c\in\mathbb{R}\setminus\{0\}$, then $x_{n}\to c$, forcing $c\in[1,2]$ (closed in $\overline{\mathbb{R}}$), so $L\in K$.
  • $L=\ast$ is impossible: by Proposition 4.2, $\hat{x}_{n}\to\ast$ requires $(x_{n})$ to have no accumulation point in $\mathbb{R}\setminus\{0\}$, but $(x_{n})\subseteq[1,2]$ is bounded and the Bolzano–Weierstrass theorem yields an accumulation point in $1,2]\subset\mathbb{R}\setminus\{0\}$.
Thus $K$ contains all its sequential limits but is not closed. ∎
Remark 4.8.
This confirms that the topology at $\ast$ is genuinely not determined by sequences: nets carry strictly more topological information than sequences (see also Example 4.5). For general background on sequential spaces, see [6] and [8].
5 ERI in the Wilansky Hierarchy
This section contains the main results of the paper.
5.1 ERI is US
Theorem 5.1.
$(X,\tau_{\mathrm{ERI}})$ is a US-space.
Proof.
Immediate from Proposition 4.4: every convergent sequence in $X$ has a unique limit. ∎
5.2 ERI is not KC
Theorem 5.2.
$(X,\tau_{\mathrm{ERI}})$ is not a KC-space.
Proof.
Let $a,b]\subset\mathbb{R}\setminus\{0\}$ with $a<b$ (e.g., $[a,b]=[1,2]$). Set $K:=q([a,b])$.
$K$ is compact.  The map $q$ restricted to $[a,b]$ is continuous (Proposition 2.12) and $[a,b]$ is compact in $\overline{\mathbb{R}}$, so $K=q([a,b])$ is compact in $X$.
$K$ is not closed.  We show $X\setminus K\notin\tau_{\mathrm{ERI}}$. Since $\{-\infty,0,+\infty\}\cap[a,b]=\varnothing$ (as $a,b]\subset\mathbb{R}\setminus\{0\}$), we have $\ast\notin K$, so $\ast\in X\setminus K$. The preimage is
$q^{-1}(X\setminus K)\;=\;\overline{\mathbb{R}}\setminus[a,b],$
which is open in $\overline{\mathbb{R}}$ (complement of a closed set). However, $\overline{\mathbb{R}}\setminus[a,b]$ is not dense: its complement $[a,b]$ has non-empty interior $(a,b)$. Therefore condition (b) fails, and $X\setminus K\notin\tau_{\mathrm{ERI}}$.
Since $K$ is compact but not closed, $X$ is not KC. ∎
Remark 5.3.
More generally, for any closed subset $F\subset\mathbb{R}\setminus\{0\}$ with $\mathrm{int}(F)\neq\varnothing$ (i.e., $F$ is not nowhere dense), the set $q(F)$ is compact but not closed in $X$. The KC failure is thus pervasive: the only compact subsets that are closed are those whose preimage complement is dense—precisely the nowhere-dense ones and those not involving $\ast$.
5.3 The role of first-countability
The following classical result (Fact 1.2 from the Introduction) explains why ERI can be US without being $T_{2}$:
Proposition 5.4 (Classical).
A first-countable space is Hausdorff if and only if it is US.
Proof.
($\Rightarrow$) is standard ([7], Th. 17.10). ($\Leftarrow$): suppose $X$ is first-countable and not Hausdorff. Let $x\neq y$ be inseparable. Let $\{U_{n}\}$ and $\{V_{n}\}$ be nested countable bases at $x$ and $y$. For each $n$, pick $z_{n}\in U_{n}\cap V_{n}$ (non-empty by inseparability). Then $z_{n}\to x$ and $z_{n}\to y$, so $X$ is not US. ∎
Corollary 5.5.
The failure of first-countability at $\ast$ (Proposition 4.1) is the precise obstruction that allows $X$ to be US without being $T_{2}$. More precisely:
  1. (i)
    At every point $\hat{x}\neq\ast$, the space $X$ is first-countable (the sets q((x1/n,x+1/n))q((x-1/n,x+1/n)) for $n>1/|x|$ form a countable base).
  2. (ii)
    If $X$ were first-countable at $\ast$, then $\mathrm{US}+\text{first-countable}$ would force $T_{2}$ by Proposition 5.4—contradicting Theorem 3.2.
  3. (iii)
    The failure of first-countability at $\ast$ is therefore necessary for the coexistence of US and non-Hausdorff.
5.4 Summary: ERI in the hierarchy
Combining the results:
Property $T_{1}$ US KC $T_{2}$
ERI $\times$ $\times$
Thus ERI provides a concrete example establishing the strict implication $\mathrm{KC}\Rightarrow\mathrm{US}$ (i.e., showing US $\not\Rightarrow$ KC) with a compact, path-connected space arising from an elementary quotient construction.
6 Comparison with Known Examples
Definition 6.1 (Reference examples).
  1. (i)
    The particular point topology (PPT) on a set $Y$ with distinguished point $p$: open sets are $\varnothing$ and all subsets containing $p$ ([9], Ex. 8–9).
  2. (ii)
    The line with two origins (LTO): the quotient of $\mathbb{R}\times\{0,1\}$ identifying $(x,0)\sim(x,1)$ for $x\neq 0$.
  3. (iii)
    The cocountable topology on an uncountable set $Y$: closed sets are countable sets and $Y$ itself.
  4. (iv)
    Van Douwen’s space [3]: a countable, compact, Fréchet–Urysohn, anti-Hausdorff US-space constructed via MAD families.
Space $T_{1}$ US KC $T_{2}$ Compact 1st-ct. Path-conn.
ERI $\times$ $\times$ $\times^{\ast}$
PPT ($|Y|{=}\infty$) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
LTO $\times$ $\times$ $\times$ $\times$
Cocountable $\times$ $\times$ $\times$ $\times$
Van Douwen $\times$ $\times^{\dagger}$ $\times$
$(X_{0},\tau_{q})$
 Fails only at $\ast$.    Fréchet–Urysohn but not first-countable; by Fact 1.2, first-countable ++ US would force $T_{2}$.    By Künzi–van der Zypen [4]: Fréchet ++ US $\Rightarrow$ KC.
Key distinctions:
  1. 1.
    ERI vs. cocountable:  Both are $T_{1}$, US, not $T_{2}$. The cocountable topology is KC but not compact; ERI is compact but not KC. In the cocountable topology, the only convergent sequences are eventually constant (the US property is “vacuous”); in ERI, convergence to $\ast$ has a precise, non-trivial criterion (Proposition 4.2).
  2. 2.
    ERI vs. van Douwen:  Both are compact and US. Van Douwen’s space is KC (and anti-Hausdorff); ERI is not KC. Van Douwen’s construction uses MAD families; ERI is completely explicit.
  3. 3.
    ERI vs. LTO:  LTO is first-countable and $T_{1}$ but not US ($1/n$ converges to both origins). ERI is US but not first-countable at $\ast$. By Proposition 5.4, this is not coincidental: first-countability ++ $T_{1}$ ++ not-$T_{2}$ forces not-US.
  4. 4.
    ERI vs. $(X_{0},\tau_{q})$:  The standard quotient is Hausdorff and first-countable. Condition (b) is the sole mechanism that destroys $T_{2}$, first-countability at $\ast$, and KC, while preserving $T_{1}$ and US.
7 Generalization
Lemma 7.1.
A Hausdorff space with no isolated points and at least one point is infinite.
Proof.
A finite Hausdorff space is discrete; in a discrete space every point is isolated. ∎
Theorem 7.2 (Generalized ERI construction).
Let $Y$ be a compact Hausdorff space with no isolated points, and let $A=\{a_{1},\ldots,a_{k}\}\subset Y$ ($k\geq 1$) be a finite non-empty subset. Let $Y_{0}=Y/A$ with projection $q:Y\to Y_{0}$ and collapsed point $\ast:=q(A)$. Define the generalized ERI topology $\tau$ on $Y_{0}$ by: $U\in\tau$ iff $q^{-1}(U)$ is open in $Y$ and $\ast\in U\Rightarrow q^{-1}(U)$ is dense in $Y$. Then $(Y_{0},\tau)$ is:
  1. (a)
    a well-defined topology;
  2. (b)
    compact;
  3. (c)
    $T_{1}$;
  4. (d)
    not Hausdorff;
  5. (e)
    connected;
  6. (f)
    US;
  7. (g)
    not KC.
Proof.
Parts (a)–(d) follow by the same arguments as for ERI (replacing $\overline{\mathbb{R}}$ with $Y$ and $\{-\infty,0,+\infty\}$ with $A$); the proofs in Sections 24 use only the compact Hausdorff and no-isolated-points properties of the ambient space and the finiteness of $A$.
(e) Connected.  Suppose $Y_{0}=U\cup V$ is a disconnection with $\ast\in U$. Then $q^{-1}(U)$ is open and dense (by condition (b), since $\ast\in U$), and $q^{-1}(V)$ is open and nonempty. But a dense set meets every nonempty open set, so $q^{-1}(U)\cap q^{-1}(V)\neq\varnothing$, contradicting $U\cap V=\varnothing$.
(f) US.  We first establish the convergence criterion in the general setting. Let $(z_{n})$ be a sequence in $Y\setminus A$ and $\hat{z}_{n}:=q(z_{n})$.
Claim: $\hat{z}_{n}\to\ast$ in $(Y_{0},\tau)$ if and only if $(z_{n})$ has no accumulation point in $Y\setminus A$.
($\Rightarrow$): Suppose $t\in Y\setminus A$ is an accumulation point. A subsequence satisfies $z_{n_{k}}\to t$ in $Y$. Let $C:=\{z_{n_{k}}:k\geq 1\}\cup\{t\}$. Since $Y$ is Hausdorff, $t$ is the only accumulation point of $C$, and $t\in C$, so $C$ is closed. $C$ is countable, hence nowhere dense: if $\mathrm{int}_{Y}(C)\neq\varnothing$, then $C$ contains a non-empty open subset of $Y$, which is uncountable in any compact Hausdorff space without isolated points (by the Baire Category Theorem)—contradicting countability. Since $C\subset Y\setminus A$, the set $q(Y\setminus C)$ is an open neighborhood of $\ast$ in $\tau$ (preimage $Y\setminus C$ is open and dense), but $\hat{z}_{n_{k}}\notin q(Y\setminus C)$ for all kk—contradicting $\hat{z}_{n}\to\ast$.
($\Leftarrow$): Let $U$ be a neighborhood of $\ast$ in $\tau$. Then $F:=Y\setminus q^{-1}(U)$ is closed in $Y$ (hence compact) and contained in $Y\setminus A$. If infinitely many $z_{n}$ belonged to $F$, compactness yields a convergent subsequence with limit in $F\subset Y\setminus A$, contradicting the hypothesis. Hence eventually $\hat{z}_{n}\in U$.
With this criterion, unique sequential limits follow by the same case analysis as Proposition 4.4. If $y_{n}=\ast$ for infinitely many $n$, the $T_{1}$ property forces $L=\ast$ as the only possible limit. Otherwise reduce to $\hat{z}_{n}$ with $z_{n}\in Y\setminus A$. If $\hat{z}_{n}\to\ast$ and $\hat{z}_{n}\to\hat{c}$ for $c\in Y\setminus A$, then $z_{n}\to c$ in $Y$ (since $q$ restricts to a homeomorphism $Y\setminus A\cong Y_{0}\setminus\{\ast\}$), making $c$ an accumulation point in $Y\setminus A$—contradicting $\hat{z}_{n}\to\ast$ by the criterion above. If $\hat{z}_{n}\to\hat{c}$ and $\hat{z}_{n}\to\hat{d}$ with $c\neq d$ both in $Y\setminus A$, then $z_{n}\to c$ and $z_{n}\to d$ in $Y$, contradicting the Hausdorff property of $Y$.
(g) Not KC.  By Lemma 7.1, $Y$ is infinite, so $Y\setminus A$ is non-empty and dense in $Y$ (complement of a finite closed set in a space without isolated points). Since $Y$ is compact Hausdorff (hence regular [8]), for any $y\in Y\setminus A$ there exists an open set $W$ with $y\in W\subseteq\mathrm{cl}(W)\subseteq Y\setminus A$. Set $C:=\mathrm{cl}(W)$. Then $C$ is closed in $Y$ with $\mathrm{int}_{Y}(C)\supseteq W\neq\varnothing$, and $C\subset Y\setminus A$. Now $q(C)$ is compact (continuous image) and $\ast\in Y_{0}\setminus q(C)$ (since $C\cap A=\varnothing$), but $q^{-1}(Y_{0}\setminus q(C))=Y\setminus C$ is not dense (since $\mathrm{int}(C)\neq\varnothing$). So condition (b) fails and $q(C)$ is not closed. ∎
Remark 7.3.
Connectedness of $Y$ is not required; the density condition forces $(Y_{0},\tau)$ to be connected even when $Y$ is disconnected (e.g., $Y=[0,1]\sqcup[2,3]$). ERI is the instance $Y=\overline{\mathbb{R}}$, $A=\{-\infty,0,+\infty\}$. Path-connectedness does not generalize without further hypotheses on $Y$.
8 The Filter-Modified Quotient Framework
The density condition in Definition 2.1 and its generalization (Theorem 7.2) are instances of a construction that can be stated for arbitrary quotient spaces and arbitrary “filter-like” requirements.
Definition 8.1 (Admissible modifier).
Let $(X,\tau)$ be a topological space. A collection $\mathcal{F}\subseteq\mathcal{P}(X)$ is a $\tau$-admissible modifier if:
  1. (F1)
    $X\in\mathcal{F}$.
  2. (F2)
    $\mathcal{F}$ is upward closed: $F\in\mathcal{F}$ and $F\subseteq G$ implies $G\in\mathcal{F}$.
  3. (F3)
    If $U_{1},U_{2}\in\mathcal{F}\cap\tau$, then $U_{1}\cap U_{2}\in\mathcal{F}$.
Definition 8.2 (Filter-Modified Quotient).
Let $(X,\tau)$ be a topological space, $A\subseteq X$ nonempty, $q\colon X\to X/A$ the projection collapsing $A$ to $\ast:=q(A)$, and $\mathcal{F}$ a $\tau$-admissible modifier. The $\mathcal{F}$-modified quotient topology is
$\tau_{\mathcal{F}}\;:=\;\bigl\{\,U\subseteq X/A:q^{-1}(U)\in\tau\;\;\text{and}\;\;\bigl(\ast\in U\Longrightarrow q^{-1}(U)\in\mathcal{F}\bigr)\,\bigr\}.$
We write $\mathrm{FMQ}(X,A,\mathcal{F}):=(X/A,\,\tau_{\mathcal{F}})$.
Proposition 8.3.
$\tau_{\mathcal{F}}$ is a topology on $X/A$, with $\tau_{\mathcal{F}}\subseteq\tau_{q}$ (the standard quotient topology) and equality on subsets not containing $\ast$. Moreover, $q\colon(X,\tau)\to(X/A,\tau_{\mathcal{F}})$ is continuous.
Proof.
$\varnothing\in\tau_{\mathcal{F}}$ (vacuous) and $X/A\in\tau_{\mathcal{F}}$ (since $X\in\mathcal{F}$ by (F1)). For finite intersections: if $\ast\in U_{1}\cap U_{2}$, then $q^{-1}(U_{i})\in\mathcal{F}\cap\tau$, and (F3) gives $q^{-1}(U_{1})\cap q^{-1}(U_{2})\in\mathcal{F}$. For arbitrary unions: if $\ast\in\bigcup_{i}U_{i}$, pick $j$ with $\ast\in U_{j}$; then $q^{-1}(U_{j})\in\mathcal{F}$ and $q^{-1}(U_{j})\subseteq\bigcup_{i}q^{-1}(U_{i})$, so (F2) gives $\bigcup_{i}q^{-1}(U_{i})\in\mathcal{F}$. The remaining assertions are immediate from the definition. ∎
Proposition 8.4.
For any topological space $(X,\tau)$, the collection of dense subsets $\mathcal{D}_{X}:=\{D\subseteq X:\overline{D}=X\}$ is a $\tau$-admissible modifier.
Proof.
(F1): $X$ is dense. (F2): supersets of dense sets are dense. (F3): if $U_{1},U_{2}$ are open and dense, then for any nonempty open $V$, $V\cap U_{1}$ is nonempty and open (density of $U_{1}$), so $(V\cap U_{1})\cap U_{2}\neq\varnothing$ (density of $U_{2}$). Hence $U_{1}\cap U_{2}$ is dense. ∎
Remark 8.5.
$\mathcal{D}_{X}$ is not a filter in general: in $\mathbb{R}$, $\mathbb{Q}$ and $\mathbb{R}\setminus\mathbb{Q}$ are both dense with empty intersection. The weaker axiom (F3)—closure under open intersections—suffices for the construction.
Remark 8.6.
The generalized ERI topology of Theorem 7.2 is $\mathrm{FMQ}(Y,A,\mathcal{D}_{Y})$. The original ERI is $\mathrm{FMQ}(\overline{\mathbb{R}},\,\{-\infty,0,+\infty\},\,\mathcal{D}_{\overline{\mathbb{R}}})$.
Remark 8.7 (Connection to Lawvere–Tierney topologies).
The density modifier $\mathcal{D}_{X}$ is the set-theoretic manifestation of the dense (or double-negation) Grothendieck topology on the frame $\mathcal{O}(X)$. In the language of topos theory, a Lawvere–Tierney topology $j\colon\Omega\to\Omega$ determines which sieves are “covers”; the dense topology declares that $U$ covers $V$ precisely when $U$ is dense in $V$. The FMQ construction with $\mathcal{F}=\mathcal{D}_{X}$ is thus a concrete point-set instantiation of sheafification with respect to the dense topology.
8.1 The mechanism: why density separates US from KC
The following lemma, used implicitly in the proofs of Proposition 4.1 and Theorem 7.2(f), is the engine behind the separation.
Lemma 8.8 (Countable Baire).
In a compact Hausdorff space without isolated points, every countable subset has empty interior.
Proof.
Let $S=\{s_{1},s_{2},\ldots\}$ be countable and suppose $\mathrm{int}(S)\neq\varnothing$; let $U\subseteq S$ be nonempty open. Since $X$ has no isolated points, neither does $U$: if $x\in U$ were isolated in $U$, then $\{x\}=U\cap V$ for some open $V$, making $\{x\}$ open in $X$. Each $U_{n}:=U\setminus\{s_{n}\}$ is open and dense in $U$ (open since $\{s_{n}\}$ is closed; dense since $s_{n}$ is not isolated). Since $X$ is compact Hausdorff it is Baire [7], and open subsets of Baire spaces are Baire, so $\bigcap_{n}U_{n}=U\setminus S$ is dense in $U$. But $U\subseteq S$ gives $U\setminus S=\varnothing$, which cannot be dense in the nonempty space $U$. ∎
The density modifier creates an asymmetry based on interior:
  • US:  a convergent sequence $\{x_{n}\}\cup\{p\}$ is countable, hence has empty interior (Lemma 8.8), hence has dense complement. That complement, projected via $q$, is an FMQ-open neighborhood of $\ast$ that the sequence eventually avoids—blocking a second limit.
  • Not KC:  a compact set $C$ with $\mathrm{int}(C)\neq\varnothing$ has non-dense complement, so $Y\setminus q(C)$ fails the density condition and is not open; $q(C)$ is compact but not closed.
8.2 The Hausdorff barrier
We now prove that, within the FMQ framework, the KC property forces Hausdorff. This establishes that the density modifier achieves the finest possible non-Hausdorff separation level.
Theorem 8.9 (Hausdorff barrier).
Let $X$ be compact Hausdorff, $A\subseteq X$ finite nonempty, and $\mathcal{F}$ any $\tau$-admissible modifier. If $\mathrm{FMQ}(X,A,\mathcal{F})$ is KC, then it is Hausdorff.
Proof.
Write $Y=\mathrm{FMQ}(X,A,\mathcal{F})$ and $\ast=q(A)$.
Separating $\ast$ from $\hat{p}$ ($p\notin A$).  Since $X$ is compact Hausdorff, it is normal. By the Urysohn lemma [7] there exists a continuous $f\colon X\to[0,1]$ with $f(p)=0$ and $f(A)=\{1\}$. Set $V:=f^{-1}\bigl([0,\tfrac{1}{2})\bigr)$; then $\overline{V}\subseteq f^{-1}\bigl([0,\tfrac{1}{2}]\bigr)$, so $\overline{V}\cap A=\varnothing$.
The set $\overline{V}$ is compact (closed in compact $X$) and $\ast\notin q(\overline{V})$ (since $\overline{V}\cap A=\varnothing$), so $q(\overline{V})$ is compact in $Y$ (continuous image of compact under $q$). By the KC hypothesis, $q(\overline{V})$ is closed, hence $U:=Y\setminus q(\overline{V})$ is open with $\ast\in U$.
Now $q(V)$ is open in $Y$ (since $\ast\notin q(V)$, the modifier condition is vacuous) and $\hat{p}\in q(V)$. Since $q(V)\subseteq q(\overline{V})$, we have $q(V)\cap U=\varnothing$. Hence $q(V)$ and $U$ are disjoint open neighborhoods of $\hat{p}$ and $\ast$.
Separating $\hat{p}_{1}$ from $\hat{p}_{2}$ with $p_{1},p_{2}\notin A$, $p_{1}\neq p_{2}$.  Since $X$ is Hausdorff, there exist disjoint open $V_{1}\ni p_{1}$, $V_{2}\ni p_{2}$ with $V_{i}\cap A=\varnothing$. Then $q(V_{1})$ and $q(V_{2})$ are disjoint open sets in $Y$ (the modifier condition is vacuous since $\ast\notin q(V_{i})$). ∎
Corollary 8.10.
For single-point-collapse FMQ spaces over compact Hausdorff base spaces, the Wilansky hierarchy collapses at KC: $\mathrm{KC}\Rightarrow T_{2}$. The density modifier $\mathcal{D}_{X}$ achieves the finest non-Hausdorff level, namely $\mathrm{US}$-not-$\mathrm{KC}$.
Remark 8.11 (Classical generalization).
Theorem 8.9 is a special case of a more general classical result: a quotient map $q\colon X\to Y$ out of a compact Hausdorff space is closed if and only if $Y$ is Hausdorff (see [8], Theorem 9.2 and the discussion therein). In a KC space every compact set is closed, so $q$ is in particular a closed map (since continuous images of compact sets in $X$ are compact in $Y$), which forces $Y$ to be Hausdorff. Thus the barrier $\mathrm{KC}\Rightarrow T_{2}$ holds for all quotients of compact Hausdorff spaces, not only for FMQ constructions. The FMQ proof via the Urysohn lemma above is more explicit and reveals the geometric mechanism: the KC hypothesis provides enough closed sets to build disjoint neighborhoods of $\ast$ and any other point.
8.3 Functional triviality
Proposition 8.12 ($C(X,\mathbb{R})$ is trivial).
Every continuous function $f\colon X\to\mathbb{R}$ is constant. That is, $C(X,\mathbb{R})\cong\mathbb{R}$.
Proof.
Let $f\colon X\to\mathbb{R}$ be continuous. Then $g:=f\circ q\colon\overline{\mathbb{R}}\to\mathbb{R}$ is continuous, and $g(-\infty)=g(0)=g(+\infty)$; call this value $c$. Suppose $f$ is not constant. Then $g$ is not constant, so there exists $t\in\mathbb{R}\setminus\{0\}$ with $g(t)\neq c$. Pick $\varepsilon<|g(t)-c|/2$. The set $U:=f^{-1}((c-\varepsilon,c+\varepsilon))$ is open in $X$ and contains $\ast$ (since $f(\ast)=c$). By the definition of $\tau_{\mathrm{ERI}}$, $q^{-1}(U)$ must be dense in $\overline{\mathbb{R}}$. But $q^{-1}(U)=g^{-1}((c-\varepsilon,c+\varepsilon))$, whose complement contains the nonempty open set $g^{-1}((g(t)-\varepsilon,g(t)+\varepsilon))$ (by continuity of $g$ and the choice of $\varepsilon$). Hence $q^{-1}(U)$ is not dense—a contradiction. ∎
Remark 8.13.
In the language of $C^{*}$-algebras, the Gelfand spectrum of $C(X,\mathbb{R})$ is a single point. This is a strong functional-analytic manifestation of the non-Hausdorff character of ERI: the topology is rich enough to support non-trivial convergence (Proposition 4.2), yet too coarse for continuous real-valued functions to detect it.
8.4 Position in the refined hierarchy
We now locate ERI in the refined hierarchy of Clontz [2]: $T_{2}\Rightarrow\mathrm{k_{1}H}\Rightarrow\mathrm{KC}\Rightarrow\mathrm{wH}\Rightarrow\mathrm{k_{2}H}\Rightarrow\mathrm{US}\Rightarrow T_{1}$, and in the chain of Bella and Costantini [11]: $\mathrm{KC}\Rightarrow\mathrm{SC}\Rightarrow\mathrm{US}$, where a space is SC (sequentially closed) if every convergent sequence together with its limit is a closed set.
Proposition 8.14.
ERI is not weakly Hausdorff.
Proof.
A space is weakly Hausdorff (wH) if every continuous image of a compact Hausdorff space is closed [2]. The map $f\colon[0,1]\to X$ defined by $f(t)=q(t+1)$ is continuous (composition of $q$ with $t\mapsto t+1$) and $[0,1]$ is compact Hausdorff. The image $f([0,1])=q([1,2])$ is not closed: $\ast\in X\setminus q([1,2])$ (since $1,2]\subset\mathbb{R}\setminus\{0\}$), but $q^{-1}(X\setminus q([1,2]))=\overline{\mathbb{R}}\setminus[1,2]$ is not dense (the complement $[1,2]$ has interior $(1,2)$), so condition (b) fails. ∎
Proposition 8.15.
ERI is SC (sequentially closed).
Proof.
Let $(y_{n})$ be a convergent sequence in $X$ with limit $L$, and set $S:=\{y_{n}:n\geq 1\}\cup\{L\}$. We show $X\setminus S\in\tau_{\mathrm{ERI}}$.
Case $L=\hat{c}$, $c\in\mathbb{R}\setminus\{0\}$.  Write $y_{n}=q(x_{n})$ with $x_{n}\to c$. Then $S\subset X\setminus\{\ast\}$ and $\ast\in X\setminus S$. The set $C:=\{c\}\cup\{x_{n}:n\geq 1\}$ is countable, closed in $\overline{\mathbb{R}}$ (limit point $c$ included), and nowhere dense. Hence $q^{-1}(X\setminus S)=\overline{\mathbb{R}}\setminus C$ is open and dense.
Case $L=\ast$.  By Proposition 4.2, $(x_{n})$ has no accumulation point in $\mathbb{R}\setminus\{0\}$; every accumulation point in $\overline{\mathbb{R}}$ belongs to $\{-\infty,0,+\infty\}$. Since $\ast\in S$, condition (b) is vacuous for $X\setminus S$. The set $\{-\infty,0,+\infty\}\cup\{x_{n}:n\geq 1\}$ is closed in $\overline{\mathbb{R}}$ (all limit points of $\{x_{n}\}$ lie in $\{-\infty,0,+\infty\}$ and are included), so $q^{-1}(X\setminus S)$ is open. ∎
Remark 8.16 (Position summary).
Combining the results of this paper:
$\underbrace{T_{2}\;\Rightarrow\;\mathrm{k_{1}H}\;\Rightarrow\;\mathrm{KC}}_{\text{ERI: No}}\;\Rightarrow\;\underbrace{\mathrm{wH}}_{\text{ERI: No}}\;\Rightarrow\;\underbrace{\mathrm{k_{2}H}}_{\text{open}}\;\Rightarrow\;\underbrace{\mathrm{US}}_{\text{ERI: Yes}}\;\Rightarrow\;\underbrace{T_{1}}_{\text{ERI: Yes}}$
ERI is also SC but not KC, witnessing the strict implication $\mathrm{KC}\Rightarrow\mathrm{SC}$. Among the known compact US-not-KC examples, ERI occupies the same Clontz level as the one-point compactification of the Arens–Fort space (pi-Base S165): both are not weakly Hausdorff. ERI is the only path-connected example at this level.
9 Discussion
9.1 The density condition as a hierarchy-calibrated mechanism
Condition (b) in the definition of $\tau_{\mathrm{ERI}}$ simultaneously:
  1. 1.
    Destroys $T_{2}$ (and KC): every neighborhood of $\ast$ has dense preimage, preventing separation from any other point and preventing compact subsets with non-nowhere-dense preimage from being closed.
  2. 2.
    Preserves $T_{1}$ (and US): the density requirement is automatically satisfied when closing singletons (by the no-isolated-points property of $\overline{\mathbb{R}}$), and the convergence criterion ensures sequential limit uniqueness.
  3. 3.
    Destroys first-countability at $\ast$: the Baire Category Theorem prevents a countable base at $\ast$, and this failure is what enables the coexistence of US and non-Hausdorff (Corollary 5.5).
9.2 Sequences vs. nets
$X$ exhibits the gap between Hausdorff and US:
  • Every convergent sequence has a unique limit (Theorem 5.1).
  • There exist nets with two limits (Example 4.5).
This gap exists precisely because first-countability fails at $\ast$: the topology at $\ast$ is not determined by sequences, so nets carry strictly more information. By Proposition 5.4, this failure is necessary: in any first-countable US space, the net-level and sequence-level separation properties coincide.
9.3 First-countability as a structural necessity
Remark 9.1.
No first-countable US space can fail to be $T_{2}$ (Proposition 5.4), regardless of compactness or path-connectedness. This confirms that ERI’s failure of first-countability at $\ast$ is not an artifact of the construction but a structural necessity for any US space that is not Hausdorff.
9.4 Open question
Is ERI $k_{2}$-Hausdorff? The remaining undetermined property in the Clontz hierarchy is the $k_{2}$-Hausdorff condition (Definition 24 of [2]). Remark 8.16 records the current status. Note that the one-point compactification of the Arens–Fort space (pi-Base S165) is $k_{2}$-Hausdorff (Lemma 28 of [2]), so a positive answer for ERI would place it at the same level as S165, while a negative answer would separate ERI strictly below S165 in the Clontz hierarchy.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author used Claude (Anthropic) in order to assist with formalizing mathematical arguments into rigorous proofs, structuring the manuscript in , and verifying claims against the existing literature. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.